# Investigating the transverse-momentum- and pseudorapidity-dependent flow vector decorrelation in p–Pb collisions with a Multi-Phase Transport model

### v1

@@ V SUMMARY
The $\mathrm{p}_{T}$- and $\eta$-dependent decorrelation of the flow vector in p–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.02 TeV are systematically studied with the AMPT model in this paper. The sensitivity of the decorrelations to the initial conditions, the partonic scatterings, the haronic rescatterings are probed by varying the configurations of the model. It is found that the observables are driven by the partonic scatterings and the initial conditions together, but has weak dependence on the hadronic rescattering. It suggests that the event-by-event fluctuations of the initial state lead to the fluctuations of the flow vector observed in the final state through the parton interactions. These results point to the possibility for employing flow vector decorrelation to investigate the transport properties of the hot and dense medium likely formed in small collision systems. Furthermore, we demonstrate that long-range jet correlations significantly influence the longitudinal decorrelation in p–Pb collisions. It provides further insights in understanding the fluctuations of the flow vector in small collision systems and has referential value for future measurements.
@@ VI Acknowledgement
The computation is completed in the HPC Platform of Wuhan Textile University. This work was supported by National Natural Science Foundation of China (No.12405158, 12147101, 12405159), the Natural Science Foundation of Hubei Province (No. 2024AFB136), the Fundamental Research Funds for the Central Universities, China University of Geosciences (Wuhan) with No.G1323523064, Key Laboratory of Quark and Lepton Physics in Central China Normal University (No. QLPL202106, QLPL2022P01, QLPL2023P01, QLPL2024P01).

### v2

@@ V SUMMARY
The decorrelation of the flow vector as well as its $\mathrm{p}_{T}$ and $\eta$ dependence in p–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.02 TeV are systematically studied with the AMPT model in this paper. The sensitivity of the decorrelations to the initial conditions, the partonic scatterings, the haronic rescatterings are probed by varying the configurations of the model. It is found that the observables are driven by the partonic scatterings and the initial conditions together, but has weak dependence on the hadronic rescattering. It suggests that the event-by-event fluctuations of the initial state lead to the fluctuations of the flow vector, then they are propagated to the observed decorrelation in final state according to the parton interactions. These results point to the possibility for employing flow vector decorrelation to investigate the transport properties of the hot and dense medium likely formed in small collision systems. In addition, we demonstrate that the contribution from the long-range jet correlations plays an important role in generating the longitudinal decorrelation in p–Pb collisions. It provides further insights in understanding the fluctuations of the flow vector in small collision systems and has referential value for future measurements.
@@ VI Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 12147101, 12405158, 12405159), the Natural Science Foundation of Hubei Province (No. 2024AFB136), the Fundamental Research Funds for the Central Universities, China University of Geosciences (Wuhan) with No.G1323523064, Key Laboratory of Quark and Lepton Physics in Central China Normal University (No. QLPL202106, QLPL2022P01, QLPL2023P01, QLPL2024P01).